OpenStudy (anonymous):

12 balloons in 2 boxes, 78 balloons in x boxes. I know the answer is x=15, but I don't know how to write a proportion for it.

OpenStudy (tylerd):

.12=2b 78=ab

OpenStudy (tylerd):

solve for b first then solve for a

OpenStudy (tylerd):

actually let me check real quick

OpenStudy (anonymous):

\[\frac{ 12 }{ 2 } = \frac{ 78 }{x}\] I usually see it like this, but tyler's way probably works too (it is a bit more complicated)

OpenStudy (anonymous):

you can type flip the fractions too! as long as everything lines up horizontally, then the ratio must hold.

OpenStudy (tylerd):

pretty sure ya

OpenStudy (anonymous):

Thank y'all so much! I see what I did wrong now, I didn't put the "1" in "12" on my calculator and kept getting 39. Thanks!

OpenStudy (triciaal):

why is the answer 15? 12:2 = 78:x x = 78/12*2 x = 13

OpenStudy (anonymous):

You are correct the answer is 13. Where did you get 15?